OpenStudy (anonymous):

Find the quadratic polynomial whose graph goes through the points (-1,4), (0,5), (2,25)

OpenStudy (anonymous):

we know it looks like \[ax^2+bx+5\] since \((0,5)\) is on the graph now you need \(a\) and \(b\)

OpenStudy (anonymous):

since \((-1,4)\) is on the graph , you know \[a(-1)^2+b(-1)+5=4\] or \[a-b=-1\]

OpenStudy (anonymous):

then the point \((2,25)\) tells you \[a(2)^2+b(2)+5=25\] or \[4a+2b=20\] solve the system \[a-b=-1\]\[4a+2b=20\] for \(a\) and \(b\)

OpenStudy (anonymous):

double the first one and add it to the second \[2a-2b=-2\]\[4a+2b=20\] \[6a=18\implies a=3\]

OpenStudy (anonymous):

and b=4?

OpenStudy (anonymous):

\[12+2b=20\] \[2b=8\implies b=4\] yes you are right

OpenStudy (anonymous):

thanks!